Melissa's mother is 20 years older than Melissa, but 3 years younger than Melissa's father. Melissa's father is 7 years younger then three times Melissa's age. How old is Melissa?

Answer by Paul(988)

Melssia'a mother is x
Melissa'a father is 3x-7
Now Messlia's mom is 3 years less than melssia's father.
Equatiion:
20+x=3x-7-3
20+x=3x-10
30=2x
x=15
Hence, Melissa is 15 years old.
Check:
3(15)-7=38
20+15=35
now 38-35=3 years younger.
Paul.